40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 389     Entry time: Wed Nov 10 22:36:59 2010
Author: tara 
Type: Summary 
Category: PMC 
Subject: current PMC's OLG TF  

The current plot for PMC's OLG TF is plotted below. The RF V is 6V, Gain slider is 14 dB.

 

The UGF is 820 Hz with phase margin (PM) = 180 - 53 = 127 degree.

 

At higher gain slider setup, the system starts to oscillate. One possible cause is the peak near 10^4 Hz which

might be the PZT's resonance frequency.

Without the notch the total gain we can increase will be limited by the peak.

I'll make a notch to damp it down.

The current spec will be ~20dB notch at 12.5 kHz, FWHM ~1kHz

 

From current setup, the optical TF should be + 16.5 dB flat, and the gain added by the gain slider is +14 dB.

previous setup we have opt TF = -5 dB and +30 dB from gain slider.

So we have improved the overall gain by ~5 dB and to UGF increased from 530 to 830 Hz.

 

Attachment 1: TF_2010_11_10.png  22 kB  | Hide | Hide all
TF_2010_11_10.png
ELOG V3.1.3-